# Retraining-Free Merging of Sparse Mixture-of-Experts via Hierarchical Clustering

### v1

@@ Abstract
Sparse Mixture-of-Experts (SMoE) models represent a significant breakthrough in large language model development. These models enable performance improvements without a proportional increase in inference costs. By selectively activating a small set of parameters during task execution, SMoEs enhance model capacity. However, their deployment remains challenging due to the substantial memory footprint required to accommodate the growing number of experts. This constraint renders them less feasible in environments with limited hardware resources. To address this challenge, we propose Hierarchical Clustering for Sparsely activated Mixture of Experts (HC-SMoE), a task-agnostic expert merging framework that reduces SMoE model parameters without retraining. Unlike previous methods, HC-SMoE employs hierarchical clustering based on expert outputs. This approach ensures that the merging process remains unaffected by routing decisions. The output-based clustering strategy captures functional similarities between experts, offering an adaptable solution for models with numerous experts. We validate our approach through extensive experiments on eight zero-shot language tasks and demonstrate its effectiveness in large-scale SMoE models such as Qwen and Mixtral. Our comprehensive results demonstrate that HC-SMoE consistently achieves strong performance, which highlights its potential for real-world deployment.
@@ 1 Introduction
The exponential growth in model parameters for Transformer-based architectures in natural language processing (NLP) has led to significant performance improvements across various tasks (Chowdhery et al., 2022; OpenAI et al., 2024; Team et al., 2024). Nevertheless, this increase in size has resulted in challenges for real-world deployment and accessibility due to heightened inference latency and computational requirements (Bommasani et al., 2022) Sparsely activated Mixture of Experts (SMoE) models have emerged as a promising solution to this challenge. SMoE architectures employ a sparse activation mechanism, wherein only a subset of the model’s parameters, or ‘experts’, are activated for each input token. This design enables extensive parametric capacity without a proportional increase in computational cost during inference, as shown in previous works (Shazeer et al., 2017; Fedus et al., 2022). However, despite improvements in inference latency, the overall size of SMoE architectures poses a significant challenge for memory usage, and the efficient reduction of SMoE model size during inference has become a critical area of concern. Recently, the authors in Liu et al. (2023) have identified high representational similarity among experts and suggested a solution to enhance expert diversity within an SMoE. Such an observation is further substantiated by the empirical results provided in Lu et al. (2024). The collective evidence from these previous studies therefore suggests that model parameters in contemporary SMoE architectures may be redundant, which points to avenues for potential optimization and efficiency improvements.
@@
To address the challenges associated with redundant parameters in SMoE models, researchers have proposed various approaches in the past few years. Early endeavors focused on task-specific expert pruning (Chen et al., 2022), which progressively eliminates non-essential experts and ultimately results in a single-expert dense model tailored for a specific downstream task. However, such approaches often necessitate extensive fine-tuning to mitigate performance degradation caused by pruning. Inspired by this limitation, several studies (Lu et al., 2024; He et al., 2024) have explored retraining-free expert pruning methods. For instance, Lu et al. (2024) proposed directly trimming experts by using the smallest output loss relative to the original model as an indicator of importance, without any fine-tuning. On the other hand, He et al. (2024) introduced a more scalable pruning process based on routing scores. The other line of studies (Li et al., 2024) suggested replacing pruning approaches with a merging method to more effectively leverage shared knowledge among experts and consolidate information from the most significant ones. Nonetheless, according to our experimental investigation in 4, this method exhibits inferior generalizability in task-agnostic setups.
Figure 1: A comparison of the average accuracy across eight LM-Harness benchmarks on Qwen1.5-MoE-A2.7B-Chat (Team, 2024). HC-SMoE (Ours) outperforms existing retraining-free pruning and merging baselines, achieving superior results while removing 25% and 50% of SMoE expert parameters.
@@
In light of the aforementioned challenges, this paper proposes a retraining-free, scalable, and task-agnostic framework, termed Hierarchical Clustering for Sparsely Activated Mixture of Experts (HC-SMoE), that concentrates on reducing the parameters of an SMoE model through hierarchical clustering based on expert outputs and frequency-weighted merging. The hierarchical clustering methodology offers two key advantages. First, instead of using the grouping approach adopted in Li et al. (2024), which only calculates the similarity among the experts once, hierarchical clustering can better maintain inter-cluster diversity and intra-cluster similarity by iterative comparison. Second, in contrast to the similarity metric used in Li et al. (2024), HC-SMoE leverages experts’ outputs rather than router logits, enhancing generalizability against dataset-specific information. This observation is supported by the experimental results presented in Section 4. To evaluate our design in a task-agnostic experimental setup, we first perform clustering and merging based on the C4 dataset (Raffel et al., 2020), and then assess accuracy across eight zero-shot language tasks (Lu et al., 2024). As shown in Fig. 1, our method achieves performance comparable to the original model (i.e., Qwen) and outperforms the best-performing baseline by 6.95% and 2.14% under the 8B and 11B parameters setups, respectively. Moreover, our results in Section 4 further demonstrate that HC-SMoE outperforms all baselines with the Mixtral 8x7B setup. Our contributions are fourfold:
  • To the best of our knowledge, this study presents the first retraining-free, task-agnostic SMoE merging strategy that scales efficiently with the number of experts.
  • We demonstrate the effectiveness of using expert outputs as the similarity metric for clustering, as compared to router logits or weights employed by prior arts.
  • We highlight the importance of clustering quality prior to merging and illustrate that the proposed hierarchical clustering provides robust and reliable results for expert grouping.
  • Our experimental results reveal that HC-SMoE consistently exhibits superior performance across various benchmarks and demonstrates its efficacy on large-scale SMoE models.
@@ 2 Background and Related Works
This section presents the essential background and related works. Section 2.1 introduces the SMoE architecture, while Section 2.2 discusses various pruning and merging methods based on SMoE.
@@ 2.1 Sparsely Activated Mixture-of-Experts (SMoE)
The SMoE model comprises multiple SMoE layers, each of which contains a set of expert neural networks and a router network. Consider an input token $x$, a set of expert neural networks $\{E_{1},E_{2},...,E_{n}\}$, and a router network $R$. The output $y$ of an SMoE layer is computed as a weighted sum of the expert network outputs, which can be expressed as Eq. (1): $y=\sum_{i=1}^{n}P_{i}(x)\cdot E_{i}(x),$ (1) $E(x)=(\sigma(xW_{\text{gate}})\odot(xW_{\text{up}}))W_{\text{down}},$ (2) where $P_{i}(x)$ represents the routing score from $R$ for the $i^{th}$ expert, and $E_{i}(x)$ denotes the output of the $i^{th}$ expert network. Building upon this formulation, the experts in both Qwen (Team, 2024) and Mixtral (Jiang et al., 2024) adopt the structure of LLaMA (Touvron et al., 2023). Specifically, the feed-forward network (FFN) within each expert consists of three linear layers that function as Eq. (2), where $\odot$ signifies element-wise multiplication, $W_{\text{up}},W_{\text{gate}}\in\mathbb{R}^{d_{h}\times d_{m}}$, and $W_{\text{down}}\in\mathbb{R}^{d_{m}\times d_{h}}$ denote the weight matrices, and $\sigma$ is the activation function, specifically Sigmoid Linear Unit (SiLU, also known as the swish function) (Elfwing et al., 2018). In practice, an efficient implementation of the routing function utilizes a top-$k$ routing strategy (Shazeer et al., 2017; Fedus et al., 2022), which selects only the top $k$ experts based on the highest logits from a linear transformation of the input. The softmax operation is then applied to the $k$ largest logits, which result in a sparse activation of experts. This procedure reduces the computational overhead by activating only a small subset of the available experts. Mathematically, this procedure can be expressed as follows:
@@
where $R(x)$ represents the routing-logits and $W_{R}$ denote a learnable parameter matrix. Such a sparsely activated MoE stucture leverage this mechanism to scale efficiently while maintaining its performance. By focusing computation on the most relevant experts for each input token, the SMoE model is able to achieve a balance between the computational efficiency and the task performance.
@@ 2.2 Expert Pruning and Merging
Table 1: A Comparison of different approaches for reducing the number of experts in SMoE. Please note that F-prune is detailed in Section 4.
@@
TSEP (Chen et al., 2022) Pruning
O-prune (Lu et al., 2024) Pruning
S-prune (He et al., 2024) Pruning
F-prune Pruning
M-SMoE (Li et al., 2024) Merging
HC-SMoE (Ours) Merging
This section discusses prior methods for reducing the number of experts in an SMoE. Table 1 offers a summary of them. We begin with the pruning strategies, and then examines the merging techniques.
Pruning strategies for SMoE models have been the focus of several recent studies. Chen et al. (2022) introduced Task-Specific Expert Pruning (TSEP), which iteratively fine-tunes the model while pruning experts to gradually reduce the number of active experts for a specific downstream task. Despite its effectiveness, the extensive fine-tuning required renders this process time-consuming and computationally expensive, limiting its practicality for large-scale models. Lu et al. (2024) proposed a method which we refer to as O-prune in this study, which directly prunes experts using a combinatorial search to minimize the Output loss in a retraining-free, task-agnostic, zero-shot setting. This approach begins by identifying the number of experts to retain in each layer. It then conducts a layer-wise search across all possible expert combinations, selecting the one that minimizes output loss compared to the original model. Nevertheless, this method directly prunes experts deemed unimportant, forfeiting the opportunity to leverage their knowledge. Furthermore, the exhaustive combinatorial search becomes computationally infeasible for models with a large number of experts. For instance, reducing 50% of the experts in a model like Qwen, which has 60 experts, necessitates enumerating approximately $C(60,30)\approx 10^{18}$ combinations per layer, rendering it impractical for large-scale applications. He et al. (2024) introduced an efficient expert trimming technique, denoted as S-prune, due to its reliance on the router Score. In this approach, each expert’s router-score $P(x)$ is accumulated globally, and only the top-scoring experts are retained, while others are pruned. This strategy offers more flexibility than that in Lu et al. (2024), as it enables different layers to retain varying numbers of experts.
Model merging techniques have emerged as a promising approach to combine the strengths of multiple models. ZipIt (Stoica et al., 2024) introduces a model merging technique that allows models with the same architecture but trained on different tasks to be merged without retraining. It utilizes pairwise feature correlation to merge features both within a single model and across different models, offering flexibility in choosing correlated features. Since expert merging can be considered a multi-model merging problem, we extend ZipIt to this context. However, its extensive feature correlation computation makes it time-consuming and less effective for expert merging scenarios. M-SMoE (Li et al., 2024) proposes a three-step pipeline for expert merging in SMoE models. It first selects dominant experts based on activation frequency to decide which experts to retain in each layer, then uses router logits $R(x)$ to group experts, followed by frequency-based merging. However, in task-agnostic settings without retraining, relying on frequency information for clustering proves ineffective. This approach faces two main issues. First, frequency varies across tasks, as shown in Appendix B, making it an unreliable indicator for deciding how many experts to retain in each layer. Second, high-frequency experts within the same layer are rarely merged, overlooking their functional similarities in the feature space. Additionally, grouping based on router information is problematic, as it depends on dataset-dependent statistics. Together, the limitations can potentially hinder the model’s ability to maintain performance over diverse tasks without access to task data.
@@ 3 Methodology
In this section, we introduce the proposed HC-SMoE. Section 3.1 first provides a formal problem definition. Section 3.2 then outlines the merging pipeline. Sections 3.2.13.2.2, and 3.2.3 further explains the similarity metric, clustering method, and merging criterion of HC-SMoE, respectively.
Figure 2: (a) The general SMoE architecture. (b) Illustration of the proposed hierarchical clustering strategy based on expert outputs. (c) Comparison of expert pruning and expert merging strategies.
@@ 3.1 Problem Definition
In this study, we address the challenge of reducing the space complexity of an SMoE model through a process we term expert merging. This process aims to consolidate the existing experts in an SMoE layer into a smaller set of new experts to reduce the model’s overall size while preserving its performance. Given an initial set of experts ${E_{1},E_{2},\ldots,E_{n}}$, the merging process involves grouping these experts into subsets and combining each subset into a new expert. Mathematically, this can be expressed as: $\{E_{i_{1}},E_{i_{2}},\ldots,E_{i_{m}}\}\rightarrow E_{1}’$ ’, $\{E_{j_{1}},E_{j_{2}},\ldots,E_{j_{k}}\}\rightarrow E_{2}’$ ’, $\cdots$, $\{E_{l_{1}},E_{l_{2}},\ldots,E_{l_{p}}\}\rightarrow E_{q}^{\prime}$ where $\{E_{i_{1}},E_{i_{2}},\ldots,E_{i_{m}}\}$, $\{E_{j_{1}},E_{j_{2}},\ldots,E_{j_{k}}\}$, and so on up to $\{E_{l_{1}},E_{l_{2}},\ldots,E_{l_{p}}\}$ are disjoint subsets of the original expert set, and $E_{1}^{\prime}$, $E_{2}^{\prime}$, …, $E_{q}^{\prime}$ are the new experts resulting from the merging process. This formulation allows for the creation of multiple new experts $\{E_{1}^{\prime},E_{2}^{\prime},\ldots,E_{q}^{\prime}\}$, each derived from a distinct group of the original experts. This approach reduces the overall number of experts while maintaining the data’s original dimensionality. In such a formulation, the router network $R$ remains unmodified. If an input token was previously assigned to an expert in a merged group, it will be routed to the corresponding new expert. For example, if a token $x$ was originally routed to any expert in the set ${E_{i_{1}},E_{i_{2}},\ldots,E_{i_{m}}}$, it is then processed by $E_{1}^{\prime}$.
@@
The expert merging problem presents several unique challenges. Unlike conventional model merging methodologies, where the entities to be merged are predetermined, expert merging requires an initial clustering of experts into groups before the merging process can be executed. This additional clustering step substantially increases the complexity of the problem, as it necessitates determining which subsets of experts can be grouped in a manner that minimizes performance loss in the model. Since training an MoE model demands substantial GPU memory, we address this problem without retraining and utilize a non-benchmark dataset to collect information for the expert merging process.
3.2 Hierarchical Clustering for Sparsely Activated Mixture of Experts
In this section, we elaborate on HC-SMoE, a methodology designed with the philosophy of being retraining-free, task-agnostic, and scalable. HC-SMoE employs a two-stage pipeline comprising clustering and merging, which are illustrated in Fig. 2 (b) and the right-hand side of Fig. 2 (c), respectively. The first clustering stage determines the optimal grouping of experts for subsequent merging. Our approach demonstrates that utilizing averaged expert outputs as the similarity metric for clustering yields superior effectiveness compared to methods relying on averaged router logits or weights. Furthermore, we emphasize the critical importance of clustering quality and illustrate that hierarchical clustering with average linkage produces robust model quality post-merging. The second stage encompasses the merging of experts within each identified cluster. Our experimental findings indicate that given effective clustering, the choice of merging method (e.g., average, frequency, or alternative merging approaches) consistently results in strong performance across all benchmarks. The above components are further explained in Subsections 3.2.13.2.2, and 3.2.3.
3.2.1 Averaged Expert Outputs as Similarity Metric
Establishing a robust and reliable expert similarity metric is crucial for HC-SMoE prior to performing clustering. HC-SMoE employs an average output strategy, which proves advantageous due to its effectiveness in capturing expert functionality. The method utilizes averaged expert outputs $\frac{1}{T}\sum_{i=1}^{T}E(x_{i})$ over a calibration dataset with $T$ tokens to effectively capture each expert’s functional behavior. The expert’s output feature vectors represent the final transformed data, which encapsulates both the input’s contextual information and the expert’s learned transformations. This allows for a direct comparison of how two experts process information, serving as a better metric than conventional alternatives such as averaged router logits or expert weights (Li et al., 2024). Router logits reflect assignment preferences influnced by local factors like specific input distributions rather than the actual function similarity between experts. The routing patterns varying significantly across tasks limits their utility as task-agnostic similarity metrics. Although expert weights contain rich parameter information, their processing is computationally expensive. Furthermore, a prior study (Liu et al., 2023) has shown that SMoE experts often exhibit high similarity in the parameter space, which diminishes their effectiveness as a distinct measure of expert functionality. In contrast, averaged expert outputs provide a direct representation of each expert’s functional behavior, which therefore enhances the efficacy of identifying functionally similar experts. Prior research (Li et al., 2016; Stoica et al., 2024) indicates that models with similar outputs are likely to perform analogous functions. This correlation highlights the efficacy of clustering based on output similarity for preserving model performance after merging, and provides the foundation for HC-SMoE’s output merging strategy.
3.2.2 Hierarchical Clustering
With a reliable expert similarity metric established, the subsequent step involves selecting an appropriate clustering method. Unlike conventional model merging, which combines predefined element groups, expert merging necessitates a more flexible approach to group experts before merging. Although several potential strategies exist, as discussed in Section 4.3, our analysis reveals that hierarchical clustering (Patel et al., 2015), as illustrated in Fig. 2 (b), is well-suited for this task. This method exhibits optimal characteristics for expert grouping. Hierarchical clustering initiates by considering each expert as an individual entity and progressively merges them based on their cluster distance. This process facilitates adaptive grouping throughout the procedure. The approach continuously recalculates distances between clusters, which ensures the grouping of functionally similar experts. As a result, hierarchical clustering is able to outperform the K-means and one-shot grouping (Li et al., 2024). Section 4.3 provides an ablation study for different clustering strategies.
Our investigation encompasses three linkage methods in hierarchical clustering: single, complete, and average linkage. In single linkage, the distance between two clusters is defined by the closest pair of elements. Complete linkage employs the maximum distance, which often results in overly compact clusters that fail to capture subtle similarities. Average linkage, which considers the average pairwise distance between all elements in the clusters, achieves an optimal balance. As demonstrated in Table 4, while both single and average linkage yield satisfactory results, average linkage consistently delivers superior performance. Therefore, in this study, the average linkage method is adopted in HC-SMoE unless stated otherwise, as it ensures high intra-cluster similarity and low inter-cluster similarity. This balance effectively preserves the characteristics of the experts during merging.

### v4

# Retraining-Free Merging of Sparse MoE via Hierarchical Clustering

@@ Abstract
Sparse Mixture-of-Experts (SMoE) models represent a significant advancement in large language model (LLM) development through their efficient parameter utilization. These models achieve substantial performance improvements at reduced inference costs. However, the deployment of SMoE models faces constraints from extensive memory requirements of expert components in resource-limited environments. To address these limitations, this paper introduces Hierarchical Clustering for Sparsely activated Mixture of Experts (HC-SMoE), a task-agnostic expert merging framework for parameter reduction without retraining. HC-SMoE introduces a novel hierarchical clustering approach based on expert outputs to ensure merging robustness independent of routing decisions. The proposed output-based clustering method enables effective capture of functional relationships between experts for large-scale architectures. We provide theoretical analysis and comprehensive evaluations across multiple zero-shot language tasks to demonstrate HC-SMoE’s effectiveness in state-of-the-art models including Qwen and Mixtral. The experimental results validate HC-SMoE’s superior performance and practical applicability for real-world deployments. Our implementation is available at https://github.com/wazenmai/HC-SMoE.
Sparse Mixture-of-Experts, Merging, Compression
@@ 1 Introduction
Transformer-based architectures in natural language processing (NLP) have demonstrated significant performance improvements across various tasks with exponential parameter growth (Chowdhery et al., 2022; OpenAI et al., 2024; Team et al., 2024). This increase in size creates substantial challenges for real-world deployment due to heightened inference latency and computational demands (Bommasani et al., 2022). Sparsely activated Mixture of Experts (SMoE) models offer a promising solution through their sparse activation mechanism, where only selected model parameters (‘experts’) activate per input token. This architecture enables extensive parametric capacity without proportional computational costs during inference (Shazeer et al., 2017; Fedus et al., 2022). However, the total size of SMoE architectures presents significant memory constraints, which positions parameter reduction as a critical research priority. Recent studies have revealed important insights about expert redundancy. Liu et al. (2023) identified high representational similarities among experts and proposed methods to enhance expert diversity. Lu et al. (2024) provided additional empirical evidence to support these findings. These studies demonstrate substantial parameter redundancy in current SMoEs and highlight opportunities for optimization.
@@
Several approaches have emerged to address redundant parameters in SMoE models. Early research introduced task-specific expert pruning (Chen et al., 2022), which progressively eliminates non-essential experts and produces a single-expert dense model for specific downstream tasks. However, such approaches often necessitate extensive fine-tuning to maintain performance levels. Recent studies have advanced toward retraining-free expert pruning methods (Lu et al., 2024; He et al., 2024). Lu et al. (2024) proposed an approach to trim experts based on output loss comparison with the original model. He et al. (2024) developed a more scalable solution through routing score-based pruning. However, the complete removal of experts and their parameters can cause irreversible loss of learned representations. An alternative research direction explores expert merging rather than pruning. Li et al. (2024) introduced a method to consolidate information from significant experts. Nevertheless, our experimental results presented in Section 4 reveal limitations in this methodology’s task-agnostic generalizability.
@@
In response to these challenges, this paper introduces the Hierarchical Clustering for Sparsely Activated Mixture of Experts (HC-SMoE), a retraining-free, scalable, and task-agnostic framework. HC-SMoE reduces SMoE model parameters through hierarchical clustering based on expert outputs and frequency-weighted merging. The hierarchical clustering methodology provides two significant advantages. First, the approach advances beyond the single-pass grouping method from Li et al. (2024) through iterative comparisons, which maintains superior inter-cluster diversity and intra-cluster similarity. Second, HC-SMoE utilizes expert outputs rather than router logits as its similarity metric, which enhances generalizability beyond dataset-specific patterns. For a fair comparison with previous work, we follow standard evaluation protocols with clustering and merging on the C4 dataset (Raffel et al., 2020) and evaluate accuracy across eight zero-shot language tasks (Lu et al., 2024). Our extensive evaluations in the supplementary material further demonstrate HC-SMoE’s effectiveness across diverse dataset domains and tasks. Our results in Fig.1 illustrates HC-SMoE’s achievement of comparable performance to the original Qwen model, with improvements of $6.95\%$ and $2.14\%$ over the strongest baseline in 88B and 1111B parameter configurations. Section 4 further validates HC-SMoE’s superior performance across all baselines on Mixtral $8\times 7$B. The primary contributions of this study are summarized as:
  • This work proposes HC-SMoE, a promising retraining-free and task-agnostic merging strategy with efficient scaling characteristics for different numbers of experts.
  • We propose expert outputs as an effective similarity metric for clustering, which offers advantages over traditional router logits or weights from prior approaches.
  • Our analysis substantiates the importance of clustering quality for merging effectiveness. The proposed hierarchical clustering method produces theoretically guaranteed and empirically validated expert groupings.
  • Our extensive experiments demonstrate HC-SMoE’s consistent superior performance across diverse benchmarks and its effectiveness on multiple SMoE models.
@@ 2.1 Sparsely Activated Mixture-of-Experts (SMoE)
The SMoE model comprises multiple SMoE layers, each of which contains a set of expert neural networks and a router network. Consider an input token $x$, a set of expert neural networks $\{E_{1},E_{2},...,E_{n}\}$, and a router network $R$. The output $y$ of an SMoE layer is computed as a weighted sum of the expert network outputs, which can be expressed as:
$y=\sum_{i=1}^{n}P_{i}(x)\cdot E_{i}(x),$ (1)
$E(x)=(\sigma(xW_{\text{gate}})\odot(xW_{\text{up}}))W_{\text{down}},$ (2)
where $P_{i}(x)$ represents the $i^{th}$ expert routing score from $R$, and $E_{i}(x)$ denotes the $i^{th}$ expert network output. This architecture extends to recent models like Qwen (Team, 2024) and Mixtral (Jiang et al., 2024), which adopt the LLaMA (Touvron et al., 2023) structure. The feed-forward network (FFN) in each expert implements three linear layers as shown in Eq. (2), where element-wise multiplication $\odot$ operates with weight matrices $W_{\text{up}},W_{\text{gate}}\in\mathbb{R}^{d_{h}\times d_{m}}$, $W_{\text{down}}\in\mathbb{R}^{d_{m}\times d_{h}}$, and Sigmoid Linear Unit (SiLU) activation function $\sigma$ (Elfwing et al., 2018). The routing implementation employs an efficient top-$k$ strategy (Shazeer et al., 2017; Fedus et al., 2022) to select experts with the highest logits from linear input transformation. A subsequent softmax operation on these $k$ largest logits enables sparse expert activation, which reduces computational overhead. This selective mechanism is formulated as follows:
@@
where $R(x)$ represents routing-logits and $W_{R}$ denotes the learnable parameter matrix. This sparsely activated architecture enables efficient scaling with preserved performance through selective computation. In turn, this mechanism allows the SMoE model to optimize computational efficiency and task performance through focused expert utilization.
Figure 1: Effectiveness of expert parameter reduction approaches on Qwen1.5-MoE-A2.7B-Chat (Team, 2024). Average accuracy across 8 LM-Harness benchmarks demonstrates HC-SMoE’s superior performance over existing retraining-free pruning and merging baselines at 25%, 37.5%, and 50% expert parameter reduction rates. $\star$ indicates the original unpruned Qwen model performance.
@@ 2.2 Expert Pruning and Merging
This section reviews existing methods for expert reduction in SMoE architectures, summarized in Table 1. Recent research has focused on various pruning strategies. Chen et al. (2022) proposed Task-Specific Expert Pruning (TSEP), which reduces active experts through iterative fine-tuning for specific downstream tasks. Although effective, this approach requires extensive computational resources and time for fine-tuning, which limits its applicability to large-scale models. Lu et al. (2024) introduced a method, which we refer to as O-prune in this study, for retraining-free and task-agnostic expert reduction in zero-shot settings. The method determines expert retention counts per layer and evaluates all possible expert combinations to select configurations that minimize output deviation from the original model. However, this approach discards potential knowledge from pruned experts. Moreover, its computational requirements become prohibitive for large expert counts. For example, a 50% reduction in Qwen’s 60 experts requires evaluation of approximately $C(60,30)\approx 10^{18}$ combinations per layer. He et al. (2024) proposed an efficient expert trimming technique, denoted as S-prune, based on router scores. This method accumulates global router-scores $P(x)$ and retains top-scoring experts, which offer enhanced flexibility over O-prune by allowing variable expert retention across layers.
Table 1: A Comparison of different approaches for reducing the number of experts in SMoE. We evaluate approaches based on their retraining-free nature, task-agnostic applicability, scalability, and strategies. Our method, HC-SMoE, is compared with TSEP (Chen et al., 2022), O-prune (Lu et al., 2024), S-prune (He et al., 2024), F-prune, and M-SMoE (Li et al., 2024). Please note that F-prune is detailed in Section 4.
@@
TSEP
Pruning
O-prune
Pruning
S-prune
Pruning
F-prune
Pruning
M-SMoE
Merging
HC-SMoE
Merging
Model merging techniques have emerged as a promising approach to combine the strengths of multiple models. ZipIt (Stoica et al., 2024) introduces a model merging technique that allows models with the same architecture but trained on different tasks to be merged without retraining. It utilizes pairwise feature correlation to merge features both within a single model and across different models, and provides flexibility in choosing correlated features. Since expert merging can be considered a multi-model merging problem, we extend ZipIt to this context. However, its extensive feature correlation computation makes it time-consuming and less effective for expert merging scenarios. M-SMoE (Li et al., 2024) proposes a three-step pipeline for expert merging in SMoE models. It first selects dominant experts based on activation frequency to decide which experts to retain in each layer, then uses router logits $R(x)$ to group experts, followed by frequency-based merging. Nevertheless, in task-agnostic settings without retraining, relying on frequency information for clustering proves ineffective in Table 2 and Table 3. This approach faces three main issues. First, frequency varies across tasks, as shown in Appendix E, making it an unreliable indicator for deciding how many experts to retain in each layer. Second, high-frequency experts within the same layer are rarely merged, overlooking their functional similarities in the feature space. Moreover, grouping based on router information can be problematic, as it depends on dataset-dependent statistics. Together, the limitations can potentially hinder the model’s ability to maintain performance over diverse tasks without access to task data.
@@ 3.1 Problem Definition
In this study, we address the challenge of reducing the space complexity of an SMoE model through a process termed expert merging. This process consolidates existing experts in an SMoE layer into a smaller set while preserving the model’s performance. Each SMoE layer initially contains $n$ experts, as defined in Section 2.1. We aim to merge these experts into $r$ clusters, where $r$ represents the target number of experts after merging. For the $i$-th cluster, denoted as $C_{i}=\{E_{0}^{i},E_{1}^{i},\dots,E_{|C_{i}|}^{i}\}$, $|C_{i}|$ represents the number of original experts assigned to this cluster. Unlike conventional model merging (Yadav et al., 2023; Xu et al., 2024) with predefined element combinations, expert merging in an SMoE necessitates a two-phase procedure due to its flexible solution space: first grouping experts into clusters, then merging within each cluster. During the merging phase, experts within each cluster combine into a single new expert, which reduces the total number of experts to $r$. The distribution of original experts across clusters satisfies $\sum_{i=1}^{r}|C_{i}|=n$, which ensures that all original experts are accounted for in the merging process.
The grouping phase encompasses two distinct strategies. Static grouping maintains exactly $r$ experts per layer after merging, while dynamic grouping permits variable expert numbers per layer with an average of $r$. HC-SMoE implements static grouping, aligning with O-prune, whereas F-prune and S-prune in Table 1 utilize dynamic grouping. The router network $R$ after expert merging remains unchanged throughout this process, with inputs previously routed to any expert within a merged group now directed to the corresponding merged expert, as illustrated in Fig. 3. This preserves router dimensionality while achieving expert reduction.
@@
The expert merging problem presents several unique challenges absent in conventional model merging. First, experts exhibit high-dimensional, task-specific functional overlap that cannot be trivially disentangled through parameter-space similarity alone. Second, the lack of predefined grouping criteria necessitates clustering methods for determining which subsets of experts can be grouped in a manner that minimizes performance loss in the model. Since training an MoE model demands substantial GPU memory, we address this problem without retraining and utilize a non-benchmark dataset to collect information for the expert merging process as calibration dataset.
3.2 Hierarchical Clustering for Sparsely Activated MoE
In light of the challenges from expert clustering and merging, we introduce a new framework for SMoE model compression through expert parameter merging. Our approach achieves the balance between model size reduction, performance retention, and computational efficiency without retraining requirements. It advances beyond previous approaches by addressing three crucial aspects of SMoE compression: (1) expert similarity metrics, (2) grouping methodology, and (3) merging strategies. We present that utilizing averaged expert outputs as similarity metrics (Section 3.2.1) preserves functional equivalence among grouped experts more effectively than router-space comparisons. We then demonstrate that our hierarchical clustering approach (Section 3.2.2) enables progressive expert grouping with reduced sensitivity to initialization variations. Finally, we elaborate on the expert merging process and demonstrate that effective clustering enables our method to preserve the capabilities of the original model across diverse merging strategies (Section 3.2.3). The integration of these components enables HC-SMoE to produce reliable and stable cluster quality, and superior in addressing compression-performance trade-off.
3.2.1 Expert Outputs as Similarity Metric
The primary objective of expert merging process is to minimize functional divergence between the compressed and original models. M-SMoE falls short in this objective as it adopts router-logits as the similarity metric, which leads to the deviation of the function outputs. Motivated by evidence that output similarity correlates with functional equivalence (Li et al., 2016; Stoica et al., 2024), we propose utilizing average expert outputs over a calibration dataset $\mathcal{D}_{cal}$ with $T$ tokens to address the issue. Specifically, for expert $E_{j}$, the representative vector computation follows:
$o_{j}:=\mathbb{E}_{x\sim\mathcal{D}_{cal}}[E_{j}(x)]=\frac{1}{T}\sum_{x\in\mathcal{D}_{cal}}^{T}E(x).\vskip-5.0pt$ (4)
This formulation captures both contextual input information and expert learned transformations, validated by L2 errors of last layer outputs presented in Table 23 in the Appendix.
Prior arts face fundamental limitations in this context. For instance, router logits $R(x)$ capture input-dependent assignment preferences rather than intrinsic expert functionality, which creates task-specific biases that hinder generalization. Parameter-space comparisons, such as concatenating flattened weights (e.g., $W_{\text{gate}},W_{\text{down}},W_{\text{up}}$), face computational inefficiency and require $O(3d^{2})$ operations for $d$-dimensional experts and parameter redundancy in SMoE structures, as shown by Liu et al. (2023). In contrast, our average-output-based metric operates in $O(d)$ space with reduced memory overhead, aligns directly with the merging objective, and maintains consistency across input distribution shifts. This property ensures robustness across diverse tasks, as validated in Section 4.3, where HC-SMoE outperforms router-logits- and weight-based baselines. Note that additional quantitative results are available in Appendix D.
Figure 2: Illustration of the proposed hierarchical clustering strategy based on expert outputs. Each blue circle denotes the outputs of an expert in the embedding space. Hierarchical clustering would iteratively group the expert clusters with minimum cluster distance.
Figure 3: Comparison of expert pruning and merging strategies.
3.2.2 Hierarchical Clustering of Experts
With a reliable expert similarity metric established, the subsequent step involves clustering SMoE experts into $r$ groups for the merging process. To achieve this objective, we employ Hierarchical Clustering (HC) as the core mechanism for grouping experts based on its capability to dynamically adapt cluster assignments while maintaining initialization robustness. Unlike static partitioning methods, HC combines experts through a bottom-up agglomerative process: starting with each expert as a singleton cluster, it recursively combines the most functionally similar pairs while continuously recalculating inter-cluster distances. This iterative recalibration reflects current functional affinities of evolving clusters and enables adaptation to emergent behaviors, a capability absent in static partitioning based approaches.
Prior approaches suffer from fundamental limitations in this context. The one-shot grouping strategy proposed by Li et al. (2024), for instance, freezes cluster assignments after an initial partitioning, disregarding how merged experts alter the functional landscape of remaining clusters. Similarly, K-means (Ikotun et al., 2023) imposes restrictive assumptions about cluster geometry (e.g., spherical, equally sized) and exhibits sensitivity to centroid initialization, a flaw amplified in high-dimensional expert output spaces where random initialization often traps clusters in suboptimal local minima (Table 5 and Appendix D). HC circumvents these issues through its dendrogram-based structure, which accommodates heterogeneous cluster shapes and sizes without requiring explicit assumptions. HC recursively applies deterministic linkage criteria (e.g., average linkage) to preserve hierarchical expert relationships while mitigating initialization dependencies. The clustering process requires two essential components: (1) a distance metric for measuring differences between expert output vectors, and (2) a linkage strategy for determining inter-cluster distances. Our implementation uses the Euclidean distance, expressed as:
[…]
where $e_{i}$ and $e_{j}$ represent the metric values for computing distances between experts $i$ and $j$. For the linkage strategy, we investigate three methods: single, complete, and average:
$\text{single: }\quad\min_{a\in A,b\in B}d(a,b),$ (6)
$\text{complete: }\quad\max_{a\in A,b\in B}d(a,b),$ (7)
$\text{average: }\quad\frac{1}{|A|\cdot|B|}\sum_{a\in A}\sum_{b\in B}d(a,b),$ (8)
where $A$ and $B$ represent clusters, and $a$ and $b$ denote experts that belong to these clusters. Single linkage defines cluster distances through the closest pair of elements, while complete linkage uses the maximum distance and often produces overly compact clusters that miss subtle similarities. Average linkage considers the mean pairwise distance between cluster elements and achieves an optimal balance. As a result, the proposed HC-SMoE framework employs average-linkage HC to optimize the trade-off between intra-cluster homogeneity and inter-cluster distinctiveness. The theoretical justification is provided in Appendix A.